OpenStudy (anonymous):

∫(dx)/[x/(x^2-1)^2],solve using partial fraction

OpenStudy (anonymous):

simple, split \[x^{2}-1 = (x+1)(x-1)\]

OpenStudy (anonymous):

y do u want to do this partial fractions jus substitute x^2 = t nd ull get ur ans. in seconds

OpenStudy (anonymous):

But myabe they are asked to solve using partial fractions!

OpenStudy (anonymous):

rather substiture x^2-1 = t will be more easy

OpenStudy (anonymous):

@mashy i still havnt studied partial fractions ... i will b completing dat till tomorow den i Can Answer this question :)

OpenStudy (anonymous):

yes manjotpal bhaiya is right mashy

OpenStudy (anonymous):

(x^2-1) whole thing squared again kushashwa

OpenStudy (anonymous):

i didn't understand what u want to say mathstina

OpenStudy (anonymous):

if u r a beginner in partial fraction then pls follow the following LINKS

OpenStudy (anonymous):

if this does not help or if this is not that u wanted then tell

OpenStudy (anonymous):

no i have books that deal to a great lot extent in Caculus... but ill folloe the above links too . thnx :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

but i was mainly saying to the asker of this question but u can also view it bhaiya

OpenStudy (anonymous):

joe, can u solve?

OpenStudy (anonymous):

im a little confused as to what the function looks like...is it this? \[\frac{dx}{\frac{x}{x^{2}-1}}\]

OpenStudy (anonymous):

or this: \[\frac{x}{x^{2}-1}dx\]

OpenStudy (anonymous):

yes,first one

OpenStudy (anonymous):

then you cant really apply partial fractions because: \[\frac{1}{\frac{x}{x^{2}-1}}dx = \frac{x^2-1}{x}dx\] then you would just divide it out to get x-(1/x), and the integral would be x^2/2-ln(x)

OpenStudy (anonymous):

the second one looks like a good candidate for partial fractions.

OpenStudy (anonymous):

yes,first one

OpenStudy (anonymous):

thanks for ur help. i need to sign out nw.